company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 203.5-cm and a standard deviation of 2.4-cm. For shipment, 7 steel rods are bundled together.

Find P94, which is the average length separating the smallest 94% bundles from the largest 6% bundles.
P94 = -cm

Enter your answer as a number accurate to 2 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution :

Let X be a random variable which represents the length of steel rods produced by a company.

Given that, X ~ N(203.5, 2.4²)

μ = 203.5 cm and  σ = 2.4 cm

Bundle size (n) = 7

We have to find P​​​​​​94.

Since, P​​​​​​94 is the average length separating the smallest 94% bundles from the largest 6% bundles therefore,

P(x̄ < P​​​​​​94) = 0.94

(Where, x̄ is the average length of a bundle.)

We know that if X ~ N(μ, σ²) then, x̄ ~ N(μ, σ²/n).

And if x̄ ~ N(μ, σ²/n) then, Z = \frac{\bar{x}-\mu}{\sigma/\sqrt{n}}\sim N(0,1)

\large \therefore P\left( \frac{\bar{x}-\mu}{\sigma/\sqrt{n}} <\frac{P_{94}-\mu}{\sigma/\sqrt{n}}\right) = 0.94

\large \therefore P\left(Z <\frac{P_{94}-203.5}{2.4/\sqrt{7}}\right) = 0.94\ \ ....................(1)

Using "qnorm" function of R we get, P(Z < 1.555) = 0.94

Comparing, P(Z < 1.555) = 0.94 and (1) we get,

\large \frac{P_{94}-203.5}{2.4/\sqrt{7}} = 1.555

\large P_{94}= 203.5 +\left ( 1.555\times\frac{2.4}{\sqrt{7}} \right )

\large P_{94}= 203.5 +1.41

\large P_{94}= 204.91\ cm

Please rate the answer. Thank you.